OpenStudy (anonymous):

the equation sin(pi/6 -x) is equal to______? a. 1/2 (sqrt3 cosx-sinx) b. 1/2 (sqrt3 cos x+sin x) c. 1/2 (cos x- sqrt3 sinx) d. 1/2 (cosx+sqrt3 sinx)

OpenStudy (anonymous):

sin(A_B)=sinAcosB -cosAsinB

OpenStudy (anonymous):

so... sin30= 1/2. and cos30= root3/2. justsub the values and solve it :)

OpenStudy (anonymous):

@blueocean18 did you get it?

OpenStudy (anonymous):

yes thanks!

OpenStudy (anonymous):

:)